OpenStudy (anonymous):

Will medal! Circle A has a diameter of 7 inches, a circumference of 21.98 inches, and an area of 38.465 square inches. The diameter of circle B is 6 inches, the circumference is 18.84 inches, and the area is 28.26 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

OpenStudy (anonymous):

Openstudy is about to go down. Don't lose hope, I, and countless other helpers will be available over on http://gamutech.com/register/

OpenStudy (anonymous):

really -_-

OpenStudy (wolf1728):

Part A) circumference = PI * diameter PI = circumference / diameter PI = 21.98 / 7 PI = 3.14 PI = 18.84/6 PI = 3.14

OpenStudy (anonymous):

i just put that in my essay?

OpenStudy (wolf1728):

Part B) Area = PI * diameter^2 / 4 PI = 4 * Area / diam^2 PI = 4 * 38.465 / 7*7 PI = 153.86 / 49 PI = 3.14 Circle B PI = 4 * Area / diam^2 PI = 4 * 28.26 / 6 * 6 PI = 113.04 / 36 PI = 3.14

OpenStudy (wolf1728):

Part C) Conclusion - PI = 3.14 and is the same for both circles and is the same for EVERY circle.

OpenStudy (anonymous):

I just put all that in my essay?

OpenStudy (wolf1728):

I don't know how much detail your instructor wants so put whatever you think is needed.